INAF LBC+MODS+LUCI queue observing, 2012 October 21-22
 
Observer:  M. Faccini, E. Sani, K. Boutsia
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: M. Edwars
Instrument: MODS and LBC
 Summary 
Heavy clouds on sky at sunset. We open dome at 1:44UT.
Observation started with seeing 1.8"
We started Gilli 18C to complete yesterday's program. Due to bad seeing conditions we started the program from the beginning to have the possibility to chose the spectra taken with the right constraint
We continue with Fritz 4B, but the seeing is getting around 2 at the beginning of the target acquisition
The humidity tend to increase (73%) and the seeing to get better at the beginning of Fritz 4B integration
Program Giallongo 27A blue channel jumping after first exposure.
At 8:57UT we close for high humidity and we stay closed for the rest of the night.
  Data Summary  
	
		
			|  Project_RUN  | 
			 Objects  | 
			 Filter  | 
			 ExpTime  | 
			 Notes  | 
		
	
	
		
			|  Gilli 18C  | 
			 2XMMJ22  | 
			 red-grating  | 
			 1h 30m  | 
			 we repeated the whole program because the seeing is hight - considered COMPLETE  | 
		
		
			|  Standard  | 
			 Feige 110  | 
			 red-grating  | 
			    | 
			    | 
		
		
			|  Fritz 4B  | 
			 GG94  | 
			 dual-grating  | 
			 2h 40m  | 
			 we took an extra exposure because of seeing being slightly out of specs - considered COMPLETE  | 
		
		
			|  Standard  | 
			 G191B2b  | 
			 dual-grating / red-grating  | 
			    | 
			    | 
		
		
			|  Giallongo 27A  | 
			 MS0015  | 
			 dual channel  | 
			 15m  | 
			 blue side jumping again - we had to close for humidity  | 
		
	
  Details 
11:57 Repeat yesteeday's MODS calibs  
-  execMODS mods.3.m2rgb.cal 
mods1b 0002-0006 second flat is saturated 
  -  execMODS mods.3.m2rgbflatONLY.cal 
mods1b 0007-0016
  -  execMODS grlamps_bLS1.0.cal 
mods1b 0017-0019
  -  execMODS grlamps_r.cal 
mods1r 0002-0004
  -  execMODS grpixflats_r.cal 
mods1r 0005-0009
  -  execMODS SlitFlatLBT5_r.cal 
mods1r 0010-0014
  -  execMODS grflats_r1.0.cal 
mods1r 0015-0019
  -  execMODS grpixflats_b.cal 
mods1b 0020-0029
 
 
01:43 The clouds are gone and we open the dome
02:07 Gilli_18C$ acqMODS 2xmmj22NEW.acq 
mods1r 0020-0022
02:23 Gilli_18C$ execMODS 2xmmj22.obs 
mods1r 0023-0026
3:33 acqMODS standard.for.2xmmj22.feige110.acq 
mods1r 0027-0028 
-  modsAlign -r mods1r.20121022.0027.fits 
  -  execMOds standard.for.2xmmj22.feige110.obs 
mods1r 0029-0031
 
 
3:53  We start Fritz program B acqMODS GG94.acq 
mods1r 0032-0034 
-  on the 0033 image there is scattered light 
 
 
4:07 execMODS GG94.obs 
mods1r 0035-0037 / mods1b 0030-0032 
-  the seeing is decreasing to 1.6 at the beginning of the target acquisition
  -  at the end of the first exposure seeing is ~1.1"-1.2"
  -  execMODS GG94.obs GG94SHORT.obs 
mods1r 0038 / mods1b 0033 
  -  We repeat the integration of one frame to have the total exposure time within the seeing required by the PI. 
 
 
7:00 acqMODS g191b2b.acq 
mods1r 0039-0040 
-  modsAlign -r mods1r.2012.10.22.0039.fits
 
 
7:09 execMODS -b 
DualGrating g191b2b.obs 
mods1r 0041-0043 / mods1b 0034-0036 
-  execMODS G191B2b.obs 
mods1r 0044 -0045/ mods1b 0037 
  -  We repeat the STD because in r0041-0043 and b0034-0036 there are few counts at the faint end of the spectra
 
 
7:34 We switch to LBC * At the beginning of the switch the seeing is 1.0-1.1
8:03 pointing Giallongo 27A 
-  dofpia , /first, /X2
  -  lbcrangebal 
 
 
8:24 starting observation MS0015.9+16_GR_d  
-  light cirrus passing frough
  -  blue image elongated 083732
 
 
8:44 blue channel jumping  
 
8:47 restarting giallongo 27A d
8:57 close for humidity 
09:12 While we are closed  - we start with MODS calibrations for the Fritz program 
-  execMODS grflats.cal 
mods1r 0046-0063 / mods1b 0038-0049
  -  execMODS grlamps.cal 
mods1r 0064-0066 / mods1b 0050-0052 repeated twice because the blue channel at some point was not saving the images 
  -  execMODS grlamps.cal 
mods1r 0067-0069 / mods1b 0053-0055
  -  execMODS grpixflats.cal 
mods1r 0070-0074 / mods1b 0056-65
 
 
11:28 humidity still over 80% TO decide to not re-open the dome for the tonight
-- 
RobertoSpeziali - 21 Oct 2012